The lengths of human pregnancies are normally distributed with a mean of 268 days and a standard deviation of 15 days. What is the probability that a pregnancy lasts at least 275 days?
Answer by stanbon(75887)

The lengths of human pregnancies are normally distributed with a mean of 268 days and a standard deviation of 15 days. What is the probability that a pregnancy lasts at least 275 days?
-------------------------
z(275) = (275-268)/15 = 7/15
P(x >= 275) = P(z >= 7/15) = normalcdf(7/15,100) = 0.3204
-----------
Cheers,
Stan H.
----------